OpenStudy (anonymous):

Find the surface area of a rectangular prism with a height of 11, a length of 15, and a width of 10

OpenStudy (anas.p):

2(lw+wh+hl) i.e. \[2 \times ( (length \times width)+(width \times height)+(length \times width)(height \times length))\] So 2*((15*10+10*11+15*11) 2*(150+110+165) 2*425=850 units